OpenStudy (anonymous):

which equation represents the line that passes through the points (6 –3) and (–4 –9). the answer choices are: y+4=3/5(x+9) y+4=5/3(x+9) y-3=3/5(x+6) y+3=3/5(x-6)

geerky42 (geerky42):

\(y-y_1=m(x-x_1)\), point is \((x_1,y_1)\) To find \(m\), use this formula: \(m = \dfrac{y_2-y_1}{x_2-x_1}\)